OpenStudy (anonymous):

Will medal and fan! what is the simplified form of 3a^4 b^-2 c3? 81a^4 c^3/b^2 81a^4/b^2 c^3 3a^4/b^2 c^3 3a^4c^3/b^2

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@mathstudent55 can somebody please help me?

OpenStudy (danjs):

\[3a^4*b^{-2}*c^3\] Did i read your given expression correctly?

OpenStudy (danjs):

recall: The rule for exponents. \[\frac{ x^{-y} }{ 1 } = \frac{ 1 }{ x^y }\] If you move a base raised to an exponent to the opposite side of a fraction, the sign on the exponent will change.

OpenStudy (anonymous):

never mind i figured it out! thanks though! Ill give you a medal though!

OpenStudy (danjs):

You can imagine that the thing has a x^0 in the denominator, and by properties of exponent addition, you can see why the sign changes...\[\frac{ x^{-y }}{ x^0 } = \frac{ x^{-y}*x^y }{ x^0*x^y } = \frac{ x^0 }{ x^{0+y} } = \frac{ 1 }{ x^y }\]

OpenStudy (danjs):

a base to the zero power is equal to 1, and the fraction was multiplied by x^y / x^y or 1.

OpenStudy (danjs):

That is why the sign on the exponent changes when you switch the side of the fraction it is on.